OpenStudy (anonymous):

Paragraphs where the second and succeeding lines of the paragraph are indented have a _____. 1)first-line indent 2)right justification 3)centered 4)hanging indent

OpenStudy (anonymous):

if a=b^2/3 and b=c^3/4, what is the relation between a and c?

OpenStudy (leoh1998):

i think the answer is 4